Calculate the density of a neutron star with a radius 1.69 x 10^4 m, assuming the mass is distributed uniformly. Treat the neutron star as a giant nucleus and consider the mass of a nucleon 1.675 x 10^-27 kg. Your answer should be in the form of N x 10^17 kg/m3. Enter only the number N with two decimal places, do not enter unit.
Answer #1

Solution)

For a neutron star, n=1.4 solar mass=1.4*2*10^30

Also,

Density= mass/volume=(1.4*2*10^30)/(4/3*3.14*(1.69*10^4)^3)

Density= 1.38*10^17 kg/m^3

===========